OpenStudy (anonymous):

Consider the function below. (Use exact numbers in your answers.) f(x)=(5x)/1+ x^2 a) Find F '(2). (b) Use the answer from part (a) to find an equation of the tangent line to the curve y = F(x) at the point (2, 2).

OpenStudy (anonymous):

really need to edit this, try the equation editor

OpenStudy (anonymous):

f(x)=\[(5x) \div 1+ x ^{2}\]

OpenStudy (anonymous):

is that better? @satellite73

OpenStudy (anonymous):

no not really, but i suppose it means \[f(x)=\frac{5x}{1+x^2}\]

OpenStudy (anonymous):

and you want \(f'(x)\) right?

OpenStudy (anonymous):

for this you have no choice but the quotient rule \[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x) = 5x, f'(x)=5,g(x)=1+x^2, g'(x)=2x\]

OpenStudy (anonymous):

f' (2)

OpenStudy (anonymous):

you are going to need \(f'(x)\) to find \(f'(2)\) right?

OpenStudy (anonymous):

f'(x)=\[5\div 2x\]

OpenStudy (anonymous):

you cannot take the derivative of the numerator and denominator separately and then put one over the other you have to use the quotient rule for this

OpenStudy (anonymous):

whats the quotient rule?

OpenStudy (anonymous):

look 6 posts up

OpenStudy (anonymous):

\[5(x ^{2}-1)\div (x ^{2}+1)^{2}\]

OpenStudy (anonymous):

@satellite73